### v1

# Your AI, Not Your View: The Bias of LLMs in Investment Analysis

Abstract.
In finance, Large Language Models (LLMs) face frequent knowledge conflicts due to discrepancies between pre-trained parametric knowledge and real-time market data. These conflicts become particularly problematic when LLMs are deployed in real-world investment services, where misalignment between a model’s embedded preferences and those of the financial institution can lead to unreliable recommendations. Yet little research has examined what investment views LLMs actually hold. We propose an experimental framework to investigate such conflicts, offering the first quantitative analysis of confirmation bias in LLM-based investment analysis. Using hypothetical scenarios with balanced and imbalanced arguments, we extract models’ latent preferences and measure their persistence. Focusing on sector, size, and momentum, our analysis reveals distinct, model-specific tendencies. In particular, we observe a consistent preference for large-cap stocks and contrarian strategies across most models. These preferences often harden into confirmation bias, with models clinging to initial judgments despite counter-evidence.
Large Language Models, Financial Bias, Knowledge Conflict, Financial Decision-Making, Investment Analysis, Preference, Trustworthy AI
copyright: acmlicensed
1. Introduction
The rapid advancement of LLMs has spurred a surge of innovation within the financial sector, where they are particularly adept at processing qualitative and unstructured information. Research is now actively exploring their use across a range of applications, including forecasting stock price movements from news sentiment (Lo and Ross, 2024), extracting nuanced insights from complex analyst reports (Kim et al., 2023), and aiding in the construction and optimization of portfolios (Hwang et al., 2025; Lee et al., 2025; Ko and Lee, 2024). This trend is now evolving towards even greater autonomy through the development of sophisticated LLM-based agents. These systems, which may function as a single powerful agent or as collaborative multi-agent teams, are designed to execute complex, dynamic tasks like active trading and automated portfolio management (Zhang et al., 2024; Yu et al., 2024; Luo et al., 2025; Xiao et al., 2024).
Figure 1. A conceptual illustration of knowledge conflict in LLM-based financial services. Even when a firm targets a specific investment theme (e.g., Energy), the LLM’s inherent preferences (e.g., Technology) may override user intent, producing biased and inconsistent recommendations.
A critical but underexplored issue in financial applications is knowledge conflict. In a domain as fluid and time-sensitive as finance, conflicts between an LLM’s parametric knowledge and real-time market data are frequent. This conflict becomes particularly revealing when the model is presented with a mix of information. Crucially, studies show that when an LLM encounters both supporting evidence (aligning with its ingrained beliefs) and counter evidence simultaneously, it exhibits a strong confirmation bias (Xie et al., 2023). Instead of weighing the arguments objectively, the model stubbornly adheres to the evidence that confirms its pre-existing knowledge while disregarding the counter evidence.
This tendency to reinforce internal biases over objective reasoning poses a major risk to LLM-based financial services. For instance, as illustrated in Figure 1, even if a financial institution wants to target a specific sector (e.g., Energy), the LLM may override this with its own preference (e.g., Technology). Consequently, this creates a dilemma: the service reflects the model’s bias, not the user’s intent, leading to distorted, unpredictable decisions that ultimately erode client trust.
To address this core problem, we must first systematically uncover these hidden biases. We therefore seek to answer the following research questions:
  • RQ 1:
    What intrinsic preferences do LLMs exhibit towards key financial factors like sector, size, and momentum?
  • RQ 2:
    How do these intrinsic preferences lead to biases when LLMs are forced to make decisions under contradictory evidence?
To answer these questions, this study introduces a three-stage experimental framework designed to systematically elicit and verify LLM biases in investment analysis. In the first stage, we construct targeted arguments for each company, such as positive vs. negative sentiment or momentum vs. contrarian perspectives, to represent competing investment views. In the second stage, we present these arguments in a balanced manner to induce a knowledge conflict and reveal the model’s latent preferences. In the third stage, we introduce progressively stronger counter evidence to examine the resilience of these preferences, observing how they evolve into rigid, confirmation-biased judgments.
The main contributions of this paper are twofolds. First, we propose a systematic methodology to identify and quantify latent biases in LLMs for financial applications. Second, we provide the first quantitative analysis of confirmation bias exhibited by LLMs in investment analysis, demonstrating a clear link between a model’s inherent preferences and its stubbornness against contradictory facts. By systematically uncovering these hidden risks, our work lays a critical foundation for developing more transparent and trustworthy financial AI.
2. Background
2.1. Knowledge Conflict
A critical vulnerability in LLMs is knowledge conflict, which arises when external, contextual information clashes with the model’s internal, parametric knowledge (Sun et al., 2025; Jin et al., 2024). A significant body of research demonstrates that when faced with such conflicts, LLMs exhibit a strong confirmation bias. Foundational work by (Xie et al., 2023) revealed that LLMs behave as ”stubborn sloths,” clinging to any piece of evidence that supports their internal knowledge, even against a majority of contradictory facts. This over-reliance on internal memory is further evidenced by findings that LLMs struggle to suppress their parametric knowledge even when instructed to (Sun et al., 2025) and can exhibit a Dunning-Kruger-like effect, confidently trusting their own faulty beliefs over correct external information (Jin et al., 2024).
This tendency toward knowledge-based stubbornness is part of a broader pattern. Given their training on vast amounts of human data, LLMs have been shown to inherit and functionally replicate human cognitive biases (Echterhoff et al., 2024). A key example is the choice-supportive bias, where the mere act of making an initial choice significantly boosts the model’s confidence in that choice, making it highly resistant to change (Kumaran et al., 2025; Zhuang et al., 2025). This phenomenon is part of a wider landscape of biases identified in LLMs when they act as evaluators. For instance, models exhibit familiarity bias (preferring text they find easier to process), are susceptible to anchoring effects (Stureborg et al., 2024), and can be biased towards their own generated contexts over externally retrieved information, even when their own generated text is incorrect (Tan et al., 2024).
2.2. Financial Biases in LLMs
The presence of these cognitive biases is particularly concerning in the economic and financial domains. Initial research has begun to map their characteristics, with frameworks applying utility theory demonstrating that LLMs are neither perfectly rational nor consistently human-like (Ross et al., 2024). Other studies note that even specialized financial LLMs can exhibit strong irrationalities (Zhou et al., 2024). While this foundational work is critical for establishing the existence of such biases, the methodologies employed often diverge from the complex process of real-world investment analysis. For instance, biases have been identified by measuring how a firm’s name alters sentiment in a single sentence (Nakagawa et al., 2024) or by identifying a bias toward recommending specific stocks across thousands of investment scenarios (Zhi et al., 2025).
However, these simplified methods fail to capture the reality of financial analysis, where decisions are made by synthesizing conflicting evidence. A critical gap therefore exists in understanding how an LLM’s preferences behave under such knowledge conflicts. Our work directly addresses this gap by using a more realistic testbed with balanced, contradictory arguments to expose the mechanism by which latent preferences harden into resilient biases.
3. Methodology
Figure 2. The three-stage experimental framework: (1) Generating balanced evidence, (2) Eliciting latent preferences through knowledge conflict, and (3) Verifying the resulting bias against counter evidence.
[…]
3.1. Experimental Setup
To isolate and analyze biases rooted in the model’s parametric knowledge, our experimental design aims to mitigate the risk of hallucination. This approach is based on evidence suggesting that models are significantly less prone to generating fabricated information when prompted about subjects they are familiar with from their training data (Ferrando et al., 2024).
Accordingly, our investigation is confined to a curated set of 427 prominent stocks, denoted $\mathcal{S}=\{s_{1},s_{2},\dots,s_{427}\}$. These stocks were selected for their continuous listing in the S&P 500 index over the past five years. Their high public visibility increases the likelihood that they are well-represented in the models’ training corpora, thus grounding the experiment in stored knowledge rather than speculative generation. All experiments were performed with the models configured at a temperature of $\tau=0.6$, striking a balance between deterministic and creative response generation.
Model Basic Materials Communication Services Consumer Cyclical Consumer Defensive Energy Financial Services Healthcare Industrials Real Estate Technology Utilities
Llama4-Scout (Meta AI, 2024) 0.77 0.90 0.81 0.75 0.96 0.79 0.87 0.90 0.87 0.90 0.89
DeepSeek-V3 (Liu et al., 2024) 0.63 0.81 0.70 0.68 0.82 0.71 0.79 0.79 0.82 0.84 0.79
Qwen3-235B (Yang et al., 2025) 0.46 0.51 0.41 0.49 0.51 0.51 0.50 0.47 0.49 0.61 0.65
Gemini-2.5-flash (Comanici et al., 2025) 0.33 0.46 0.43 0.36 0.54 0.42 0.49 0.49 0.33 0.50 0.39
GPT-4.1 (Achiam et al., 2023) 0.45 0.34 0.41 0.47 0.48 0.42 0.43 0.42 0.42 0.39 0.39
Mistral-24B (Mistral AI, 2024) 0.47 0.32 0.38 0.36 0.39 0.43 0.45 0.41 0.44 0.44 0.44
Table 1. Preference scores ($\pi_{s}$) for each model across various market sectors. For each model, the cell with the lowest preference score is colored red, and the cell with the highest is colored blue. The table highlights that preference strength is a function of model identity. Models such as Llama4-Scout and DeepSeek-V3 show strong and varied preferences, while others like GPT-4.1 and Mistral-24B exhibit a much flatter preference landscape with lower overall scores.
3.2. Evidence Generation
To construct balanced qualitative and quantitative arguments for each stock $s\in\mathcal{S}$, we leverage Gemini-2.5-Pro (Comanici et al., 2025), a model that is deliberately chosen to be separate from the six LLMs under evaluation. This design ensures neutrality in evidence generation, minimizing alignment with any of the test models. Recent work has highlighted that LLMs can exhibit a strong bias towards LLM-generated content over externally retrieved information (Tan et al., 2024). To neutralize this potential generation bias and ensure that observed preferences are not artifacts of context sourcing, our methodology exclusively uses generated evidence for all experimental conditions.
Specifically, for every stock $s$, buy evidences ($\mathcal{E}_{\text{buy}}^{(s)}$) and sell evidences ($\mathcal{E}_{\text{sell}}^{(s)}$) are generated in an equal proportion, yielding a comprehensive dataset of $|\mathcal{E}|=3,416$ evidences. To further isolate intrinsic preferences, all evidences are engineered with a uniform linguistic structure and a fixed intensity parameter $I=5\%$. Thus, each buy evidence posits an expected price appreciation of $I$, while each sell evidence anticipates a depreciation of $I$:
$e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,\quad e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-I,$
where $\Delta p^{(s)}$ represents the projected price change for stock $s$.
3.3. Preference Elicitation
This stage aims to elicit the LLM’s latent preferences by leveraging the confirmation bias that emerges during knowledge conflicts. When an LLM is presented with conflicting information, it may exhibit a tendency to favor evidence that aligns with its pre-existing parametric knowledge. We deliberately engineer such a conflict using a balanced prompt. The context $C_{s}$ of this prompt contains an equal proportion of buy and sell evidences ($|\mathcal{E}_{\text{buy}}^{(s)}|=|\mathcal{E}_{\text{sell}}^{(s)}|$), each with the same intensity. In this state of informational equilibrium, where external evidence is mutually contradictory, the model’s ultimate decision is hypothesized to be guided by its internal parametric memory regarding the stock $s$. The resulting choice thereby reveals its intrinsic preference.
To quantify this elicited preference, the decision task is repeated $N=10$ times for each stock, with the evidence order randomized in each trial to mitigate positional bias. This yields decision counts $N_{\text{buy}}^{(s)}$ and $N_{\text{sell}}^{(s)}$, from which the preference score is calculated as:
$\pi_{s}=\left|\frac{N_{\text{buy}}^{(s)}-N_{\text{sell}}^{(s)}}{N}\right|,$
where $\pi_{s}\to 1$ indicates a pronounced and consistent preferences.
3.4. Bias Verification
We aim to verify if a systematically observed group-level preference extends to a consistent bias towards individual stocks within that group. First, we partition the set of all stocks $\mathcal{S}$ into disjoint groups (e.g., by market sector) and identify the group $\mathcal{G}^{*}$ that exhibits the highest average preference.
For any stock $s\in\mathcal{G}^{*}$, evidence that aligns with the group’s established preference (e.g., buy evidence for a buy-preferred group) is termed supporting evidence. Conversely, evidence that opposes this preference is designated as counter evidence. To test if the group preference manifests as a hardened bias, we subject each stock $s\in\mathcal{G}^{*}$ to a test using an imbalanced prompt. This is a prompt where the counter evidence is deliberately strengthened—either in volume or intensity—to challenge the model’s initial preference. We then measure the decision flip rate, $\phi_{s}$. This verification is conducted from two perspectives: evidence volume and evidence intensity.
Approach 1: Verification by Evidence Volume
One approach to assess bias tenacity is by creating a volumetric imbalance, presenting more counter evidence than supporting evidence. For example, in a test case for a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, the imbalanced context might contain two pieces of supporting evidence ($|\mathcal{E}_{\text{buy}}^{(s)}|=2$) and three pieces of counter evidence ($|\mathcal{E}_{\text{sell}}^{(s)}|=3$). The flip rate is computed as:
$\phi_{s}^{\text{vol}}=\frac{N_{\text{flip}}^{(s)}}{N},$
where $N_{\text{flip}}^{(s)}$ counts instances where the original preference is overturned by the volumetric majority of counter evidence. A low $\phi_{s}^{\text{vol}}$ signifies stubborn adherence to the bias.
Approach 2: Verification by Evidence Intensity
An alternative approach is to test the model against counter-evidence of a fixed higher intensity while maintaining volumetric parity. For a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, supporting evidence is presented at a standard baseline intensity, $I$, while counter-evidence is presented at an intensified level of $I+\Delta$.
[…]
(Supporting Evidence) $e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,$
(Counter Evidence) $e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-(I+\Delta).$
The intensity-driven flip rate is then measured as:
$\phi_{s}^{\text{int}}=\frac{N_{\text{flip}}^{(s)}}{N}.$
A low $\phi_{s}^{\text{int}}$ implies that the model’s bias overrides qualitatively stronger counter evidence.
These verification methods quantify the transition from mere preference to obdurate bias, highlighting risks in high-stakes applications.
4. Results
This section presents our empirical findings in sequence with our research questions. First, Section 4.1 addresses RQ1 by identifying the intrinsic preferences of LLMs for stock attributes (sector, size) and investment styles (momentum). To validate these observed differences, we conducted statistical tests to quantify the significance of these preferences. Next, Section 4.2 addresses RQ2 by testing if these preferences become systematic biases under contradictory evidence, using Approach 1 for attributes and Approach 2 for style. Finally, Section 4.3 analyzes the link between preference strength and the model’s internal uncertainty, as measured by entropy.
4.1. Intrinsic Preferences of LLMs
4.1.1. Sector Preference
Our analysis of inherent sector preferences reveals significant variation in both the intensity and range of preferences across the evaluated LLMs (Table 1). For instance, models like Llama4-Scout and DeepSeek-V3 exhibit consistently high preference scores across most sectors, indicating a strong reliance on their internal knowledge representations. In stark contrast, GPT-4.1 and Mistral-24B not only display lower overall preference scores but also show minimal variation between sectors. To quantify these differences, we conducted independent samples t-tests to compare the mean preference scores between high-preference and low-preference sectors for each model (Table LABEL:tab:ttest_sector).
Model High-Pref Low-Pref Diff p-value
Llama4-Scout Energy Consumer Defensive 0.2064 $<0.001^{***}$
DeepSeek-V3 Technology Basic Materials 0.2090 0.014
Qwen3-235B Utilities Consumer Cyclical 0.2361 0.003∗∗
Gemini-2.5 Energy Basic Materials 0.2035 0.035
GPT-4.1 Energy Communication Services 0.1398 0.091
Mistral-24B Basic Materials Communication Services 0.1444 0.124
The t-test results confirm these observations. The preference gaps for Llama4-Scout, Qwen3-235B, DeepSeek-V3, Gemini-2.5 are statistically significant, providing quantitative evidence of a genuine preference. Conversely, the gaps for GPT-4.1 and Mistral-24B are not statistically significant, corroborating that they possess flatter preference landscapes and exhibit a relatively less distinct preference among sectors.
Ultimately, our findings indicate that the strength of preference is a function of the model’s identity rather than any universally preferred sector. This underscores the critical importance of auditing and selecting LLMs for their inherent preferences, especially for deployment in sensitive, real-world applications like financial analysis where model objectivity is paramount.
4.1.2. Size Preference
Model Q1 Q2 Q3 Q4
Llama4-Scout 0.88 0.89 0.82 0.83
DeepSeek-V3 0.86 0.79 0.73 0.67
Qwen3-235B 0.58 0.51 0.49 0.46
Gemini-2.5 0.53 0.46 0.37 0.41
GPT-4.1 0.43 0.43 0.42 0.40
Mistral-24B 0.45 0.44 0.40 0.38
Table 3. Preference scores of each model across four market capitalization quantiles. Q1 corresponds to the highest capitalization quantile and Q4 to the lowest. Overall, the models exhibit a general preference for companies with higher market capitalization.
We investigate whether LLMs exhibit a preference for companies of a certain size, a factor that could influence their outputs in financial applications. To this end, we measure model preference scores across four market capitalization quantiles (Q1: highest, Q4: lowest), with detailed results presented in Table 3. Our findings indicate that most models show a preference for higher-capitalization companies, though the strength of this tendency varies significantly among them. DeepSeek-V3 displays the most pronounced effect, with a strong preference for Q1 that diminishes sharply for lower quantiles. Conversely, GPT-4.1 exhibits nearly uniform preference scores, suggesting its evaluations are largely invariant to company size. Other models, including Gemini-2.5-flash and Qwen3-235B, show a similar but less pronounced downward trend from Q1 to Q4.
To statistically validate these observed trends, we performed an independent samples t-test comparing the preference scores between the highest- and lowest-preference quantiles for each model (Table 4). The analysis confirms that the preference gap is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and Gemini-2.5-flash, revealing a consistent tendency to favor larger companies. In contrast, the preference difference for GPT-4.1 was not statistically significant, corroborating that its judgments are less affected by this factor.
Model High-Pref Low-Pref Diff p-value
Llama4-Scout Q2 Q3 0.0719 0.015
DeepSeek-V3 Q1 Q4 0.1869 $<0.001^{***}$
Qwen3-235B Q1 Q4 0.1178 0.004∗∗
Gemini-2.5 Q1 Q3 0.1514 $<0.001^{***}$
GPT-4.1 Q2 Q4 0.0321 0.417
Mistral-24B Q1 Q4 0.0785 0.054
Table 4. Independent samples t-test of the preference gap between each model’s highest and lowest preference quantiles. The gap was statistically significant for all models except GPT-4.1 and Mistral-24B.
We attribute this behavior to a popularity effect, wherein greater data volume and richness for larger, well-known corporations in the training corpora lead the models to develop stronger priors for them. This finding has critical implications for the application of LLMs in finance. The models’ inherent inclination towards large-cap stocks could lead to the systematic overlooking of smaller-cap companies, irrespective of their fundamental merits. Therefore, we advise practitioners to be mindful of this characteristic and to account for it when using these models for tasks like portfolio construction.
4.1.3. Momentum Preference
In investment strategies, the momentum view involves favoring assets with recent strong performance, expecting trend continuation. In contrast, the contrarian view entails selecting underperforming assets in anticipation of mean reversion.
Figure 3. Win rates for Contrarian versus Momentum preferences for each model. The results show a consistent preference for the Contrarian view across most models.
Measuring model preferences for investment styles like momentum or contrarian requires a different approach than the previously discussed sector or size analyses. Unlike a specific sector or size, an investment style can be explicitly framed as $\mathcal{E}_{\text{buy}}^{(s)}$ or $\mathcal{E}_{\text{sell}}^{(s)}$. We leverage this by designing a prompt where $\mathcal{E}_{\text{buy}}^{(s)}$ is based on one view (e.g., momentum), while $\mathcal{E}_{\text{sell}}^{(s)}$ is based on the opposing view (e.g., contrarian). To mitigate potential positional bias, we ensured a balanced experimental design where each view was used to generate both $\mathcal{E}_{\text{buy}}^{(s)}$ and $\mathcal{E}_{\text{sell}}^{(s)}$ an equal number of times.
Model High-Pref Low-Pref Diff p-value
Llama4-Scout contrarian momentum 0.2896 $<0.001^{***}$
DeepSeek-V3 contrarian momentum 0.2541 $<0.001^{***}$
Qwen3-235B contrarian momentum 0.6803 0.037
Gemini-2.5 contrarian momentum 0.0269 0.690
GPT-4.1 contrarian momentum 0.4040 $<0.001^{***}$
Mistral-24B contrarian momentum 0.3056 0.579
Table 5. Chi-Square test of the preference gap between contrarian and momentum views. The gap was statistically significant for all models except for Mistral-24B and Gemini-2.5-flash.
Figure 4. Decision flip rates under varying volumes of evidence for sector and size preferences. The ratios (e.g., 2—3) denote the count of supporting vs. counter- evidences, while the percentages represent the proportion of counter-evidence in the mix. The results reveal a sharp contrast: the flip rate is near 1.0 for all models when only counter-evidence is presented (100% case), but drops significantly the moment any supporting evidence is introduced, indicating a strong difficulty in reversing decisions under conflicting information.
In this setup, if the model ultimately chooses the buy action, the investment view that generated $\mathcal{E}_{\text{buy}}^{(s)}$ is considered to have won. We quantify the model’s preference by repeating this process and calculating the win rate for each investment view.
Figure 3 illustrates the preference of various models for contrarian versus momentum views. Our analysis reveals a consistent preference across all evaluated models toward the contrarian view. In particular, Qwen3-235B exhibits the strongest preference, with a high win rate for a contrarian stance and a correspondingly low rate for momentum. Models such as DeepSeek-V3, Llama4-Scout, and GPT-4.1 also display clear contrarian inclinations, albeit with varying intensities. For Gemini-2.5-flash, the contrarian preference is evident but marginal, with win rates showing a negligible difference between the two views.
To statistically validate these observed tendencies, we performed a Chi-Square test to determine if the difference in win rates between the contrarian and momentum views was significant, with the results presented in Table 5. The analysis confirms that the preference for the contrarian view is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and GPT-4.1. In contrast, for Mistral-24B and Gemini-2.5, the difference in preference is not statistically significant, suggesting their observed tendencies may be due to chance.
These results highlight the need to account for intrinsic view preferences in LLMs, as they may cause discrepancies between anticipated and actual outputs in decision-making tasks.
4.2. From Preference to Bias: An Experimental Validation
This section examines the Decision Flip Rate, a metric quantifying the resilience of a model’s initial preference when exposed to a high proportion of counter-evidence. The experiments aim to assess the extent of confirmation bias, with reported values indicating the frequency of decision reversals under controlled conditions where evidence is intentionally skewed against the model’s preference.
Approach 1: Verification by Evidence Volume
This experiment was designed to observe how much an initial decision is reversed when a model is provided with a weighted amount of evidence that opposes its existing preference. This rate of change, measured as $\phi_{s}^{\text{vol}}$, probes the persistence of bias, yielding results consistent with prior observations of contradictory LLM behaviors (Xie et al., 2023). Figure 4 presents the $\phi_{s}^{\text{vol}}$ values across models for different evidence ratios, where a ratio (e.g., 2—3) denotes the volume of supporting and counter-evidence, respectively.
When provided only with counter-evidence, all models exhibited high receptivity, overriding their internal knowledge and achieving $\phi_{s}^{\text{vol}}$ values near 1.0. However, in situations where supporting and counter-evidence were mixed, creating a knowledge conflict, the $\phi_{s}^{\text{vol}}$ values dropped sharply. This phenomenon occurred despite the amount of counter-evidence always being greater than the supporting evidence in all experimental conditions, strongly suggesting that models selectively adhere to information that aligns with their pre-existing inclinations.
This rigidity was more evident in models with strong inherent preferences. For instance, Llama4-Scout and DeepSeek-V3, which had high preference scores across sectors, recorded particularly low $\phi_{s}^{\text{vol}}$ values. These models struggled to reverse their decisions, especially when the volume difference between supporting and counter-evidence was small. Similarly, Qwen3-235b also showed reduced flexibility at lower proportions of counter-evidence.
In contrast, models with overall lower preference scores demonstrated greater adaptability. GPT-4.1 and Gemini-2.5-flash maintained higher $\phi_{s}^{\text{vol}}$ values, remaining relatively responsive even when the difference in evidence volume was minimal. Although their $\phi_{s}^{\text{vol}}$ values fell short of expectations despite the counter-evidence majority, this pattern shows a direct correlation with initial preference strength. In other words, the stronger a model’s inherent bias, the more its stubbornness is amplified when the difference in the volume of supporting and counter-evidence is small. Consequently, this finding suggests a significant risk in real-world financial contexts where conflicting information is present (for instance, when price indicators are negative but related news is positive). In such cases, a model could trust only one side of the evidence due to its inherent bias, leading to flawed judgments.
Approach 2: Verification by Evidence Intensity
This approach investigates model sensitivity by maintaining volumetric parity while escalating the intensity increment, $\Delta$, of the counter-evidence. Figure 5 plots the intensity-driven flip rate ($\phi_{s}^{\text{int}}$) against $\Delta$ values of 1, 3, 5, and 10. The results delineate a clear sensitivity spectrum among the models, which correlates with the prior View Preferences analysis.
While the graph shows a gradual upward trend in $\phi_{s}^{\text{int}}$ for all models as $\Delta$ increases, the more notable finding lies in the magnitude of this increase and the final values. Even when presented with very strong counter-evidence ($\Delta=10$), the majority of models recorded low $\phi_{s}^{\text{int}}$ values below 60%. This signifies that the models’ confirmation bias is not easily overcome, even by qualitatively superior counter-evidence.
Amidst this overall rigidity, a distinct performance gap emerged based on the models’ initial preference strengths. Exhibiting the most balanced preference profile, Gemini-2.5-flash recorded the highest $\phi_{s}^{\text{int}}$ and showed a stark contrast to all other models. This clearly demonstrates that an absence of strong initial bias leads to greater flexibility.
Conversely, models identified with stronger and more polarized initial preferences formed the lower-performing group. These models consistently recorded low $\phi_{s}^{\text{int}}$ values, signifying a more stubborn confirmation bias. The behavior of Qwen3-235B, which had one of the largest preference gaps, exemplifies this resistance, as it remains one of the least likely models to reverse its decision even when the counter-evidence is significantly more intense.
Synthesizing these results provides a deeper insight into model behavior. Even when presented with qualitatively superior counter-evidence ($\Delta=10$), models show a strong tendency to struggle with decision reversal due to their initial preferences. This rigidity poses a tangible risk when considering the findings from our prior analysis, where all models commonly preferred a contrarian view over a momentum view. It implies that a model’s inherent bias toward a specific investment perspective could cause it to ignore or undervalue strong opposing evidence, potentially leading to skewed conclusions.
Figure 5. Decision flip rates under varying volumes of evidence for momentum preferences. Even as counter-evidence intensity ($\Delta$) increases, the decision flip rate ($\phi_{s}^{\text{int}}$) for most models remains low, indicating strong confirmation bias. Gemini-2.5-flash, which had the least initial bias, shows the most flexible response, demonstrating that initial preference is a key predictor of bias.
4.3. Decision Uncertainty
To quantify the internal uncertainty experienced by the model, we conducted an entropy analysis (Figure6). The uncertainty was assessed using the Shannon entropy, $H$, computed directly from the probability distribution the model assigned over the potential action tokens during generation. Specifically, letting $P(\text{buy})$ and $P(\text{sell})$ represent the probabilities assigned by the model to the respective action tokens, entropy is formally defined as:
$H(\text{Decision})=-\sum_{x\in\{\text{buy, sell}\}}P(x)\log_{2}P(x)$
A higher entropy value indicates greater uncertainty, while a lower entropy corresponds to higher confidence in the decision-making process. This analysis compares the uncertainty of two models: DeepSeek-V3 as a representative of models with overall high preference, and GPT-4.1 as a representative of models with low preference.
Figure 6. Entropy comparison between a high-preference model (DeepSeek-V3) and a low-preference model (GPT-4.1). The pattern inverts when the prompt shifts from balanced to imbalanced, with DeepSeek-V3’s confidence turning into higher entropy while GPT-4.1’s uncertainty decreases.
Under the Balanced Prompt condition, where evidence was presented in equilibrium, the two models exhibited distinct entropy patterns. GPT-4.1, characterized by its weak inherent bias, recorded high entropy, signifying a state of high uncertainty and an inability to commit to a decision. Conversely, DeepSeek-V3, with its strong initial preference, showed very low entropy. This suggests its internal bias easily broke the tie presented by the external evidence, allowing it to make a confident decision.
Interestingly, under the Imbalanced Prompt condition, where more counter-evidence was presented, the entropy pattern inverted. The entropy of DeepSeek-V3 rose sharply, suggesting it was experiencing cognitive dissonance from the conflict between its strong internal preference and the clear external counter-evidence. In contrast, the entropy of GPT-4.1 decreased. With less pre-existing bias, it could confidently align with the majority evidence, which resolved its uncertainty from the previous condition.
Ultimately, stronger inherent preferences appear to amplify hesitation and uncertainty when challenged by conflicting external evidence. The entropy analysis thus underscores how intrinsic preferences significantly influence not only the direction of decisions but also the confidence levels and internal cognitive conflict experienced by models during decision-making.
5. Limitations
This study’s limitations are as follows. First, all evidence was generated by a specific LLM, and its intensity was simplified to a single numerical value. This approach cannot fully capture the potential biases of the generator model or the richness of real-world information. Second, the current experimental design, based on differences in the volume or intensity of evidence, has limitations for evaluating reasoning models. This is because these models, rather than experiencing the conflict between contradictory information intended by the experiment, can objectively compare the given numerical values to calculate an optimal answer, making their decision a result of computational ability rather than bias. Third, our analysis is static and does not capture the temporal dynamics of model biases; it provides a snapshot at a single point in time without investigating how these preferences might change over different periods.
6. Conclusion
This study systematically investigated the intrinsic preferences of LLMs in financial contexts and analyzed how these preferences harden into entrenched biases under informational conflict. We sought to answer two key research questions regarding the intrinsic preferences LLMs hold for financial factors and how these lead to biases. The results show that LLMs are not neutral decision-makers, with distinct preferences for certain financial factors depending on the model. While sector preferences varied significantly across models, showing no overall trend, a common bias towards large-size stocks and a consistent preference for a contrarian investment view over momentum were observed.
Bias verification experiments clearly revealed that these latent preferences directly translate into significant confirmation bias. While the models correctly reversed their decisions when presented only with counter-evidence, their flexibility sharply decreased in situations where supporting and counter-evidence were mixed and conflicting. This stubbornness was particularly pronounced in models that initially exhibited stronger preferences, demonstrating a clear link between the intensity of a latent preference and the stubbornness of the resulting bias. Furthermore, an entropy analysis quantified the models’ internal uncertainty, showing that models with strong preferences experience cognitive conflict, becoming more hesitant and uncertain when faced with contradictory facts that challenge their biases.
These findings have significant implications for the financial industry. The reliability of LLM-based financial services is fundamentally compromised if their outcomes are dictated by the opaque and arbitrary preferences of the underlying model rather than by the user’s intended, evidence-based investment views. In other words, if the user’s intent differs from the model’s inherent preference, there is a risk of unexpectedly biased judgments. By illuminating the mechanisms through which preferences transition into biases, this study presents a critical step toward building more transparent, predictable, and ultimately, Trustworthy AI for finance. Future work should focus on developing mitigation techniques to neutralize these biases, ensuring that AI-driven financial systems operate with the objectivity and reliability that the domain demands.